i) Find all points on the curve x(x + y2) = y where the tangent line is parallel to the x-axis.

ii) Find all points on the curve x(x + y2) = y where the tangent line is parallel to the y-axis
1
Expert's answer
2021-02-26T04:53:38-0500

i) Tangent line:

"(y-y_0)=y'(x_0)(x-x_0)"


"y'=x+y^2+x(1+2yy')"

"y'=\frac{x+y^2+x}{1-2xy}"


Line is parallel to the x-axis:

"y=c=const\implies y'=0"


"y'=\frac{x+y^2+x}{1-2xy}=0"


Tangent line is parallel to the x-axis if :

"2x+y^2=0"

"2xy\neq1"


ii) Tangent line is parallel to the y-axis if :

"y'\to \infin"

So we have:

"2xy=1"